If 0.42 mole of helium occupies 9.9 L, how many moles must be added to inflate the balloon to 17.0 L? (round to two sig figs)


1
Expert's answer
2021-08-20T01:52:08-0400

n1 = 0.42 mol

V1 = 9.9 L

n2 = unknown

V2 = 17.0 L


Solution:

PV = nRT

V/n = RT/P

RT/P = const

Therefore,

V1/n1 = V2/n2

n2 = n1V2 / V1

n2 = (0.42 mol × 17.0 L) / (9.9 L) = 0.72 mol

n2 = 0.72 mol


moles added:

0.72 mol - 0.42 mol = 0.30 mol


Answer: 0.30 mol of helium must be added to inflate the balloon to 15.6 L.